OpenStudy (rissaunnie):

Simplify this expression. 2x3(5x3 - 7)

OpenStudy (anonymous):

is that \[2x ^{3}\]

OpenStudy (rissaunnie):

yes

OpenStudy (anonymous):

2x^3(5x^3−7)= 2∗5x^3+3−2∗7x^3= 10x^6−14x^3

OpenStudy (anonymous):

Oh yeah, in that 2∗5x^3+3−2∗7x^3 part, I should've made it 2∗5x^(3+3)−2∗7x^3

OpenStudy (anonymous):

2x3(5x3-7) Final result : 2x3 • (5x3 - 7) Reformatting the input : Changes made to your input should not affect the solution: (1): "x3" was replaced by "x^3". 1 more similar replacement(s). Step by step solution : Step 1 : Raise x to the 3rd power Exponentiation : Equation at the end of step 1 : (2 • (x3)) • ((5 • x3) - 7) Step 2 : Raise x to the 3rd power Exponentiation : Equation at the end of step 2 : (2 • x3) • (5x3 - 7) Step 3 : Multiply 2x3 by 5x3-7 Trying to factor as a Difference of Cubes: 3.1 Factoring: 5x3-7 Theory : A difference of two perfect cubes, a3 - b3 can be factored into (a-b) • (a2 +ab +b2) Proof : (a-b)•(a2+ab+b2) = a3+a2b+ab2-ba2-b2a-b3 = a3+(a2b-ba2)+(ab2-b2a)-b3 = a3+0+0+b3 = a3+b3 Check : 5 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Polynomial Roots Calculator : 3.2 Find roots (zeroes) of : F(x) = 5x3-7 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 5 and the Trailing Constant is -7. The factor(s) are: of the Leading Coefficient : 1,5 of the Trailing Constant : 1 ,7 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 -12.00 -1 5 -0.20 -7.04 -7 1 -7.00 -1722.00 -7 5 -1.40 -20.72 1 1 1.00 -2.00 1 5 0.20 -6.96 7 1 7.00 1708.00 7 5 1.40 6.72 Polynomial Roots Calculator found no rational roots Trying to factor as a Difference of Cubes: 3.3 Factoring: 5x3-7 Check : 5 is not a cube !! Ruling : Binomial can not be factored as the difference of two perfect cubes Final result : 2x3 • (5x3 - 7)